# HAPNet: Toward Superior RGB-Thermal Scene Parsing via Hybrid, Asymmetric, and Progressive Heterogeneous Feature Fusion

### v2

@@ IV-B Evaluation Metrics
We employ two widely used evaluation metrics: accuracy (Acc) and intersection over union (IoU), to quantify the scene parsing performance with respect to each category. The mean values of these two metrics across all categories (denoted as mIoU and mAcc, respectively) are also computed to quantify the comprehensive network performance.
IV-C Implementation Details
@@
Our model was trained for 200 epochs on an NVIDIA RTX 3090 GPU, utilizing the AdamW optimizer [60] with an initial learning rate of $2\times 10^{-4}$ and a weight decay of $5\times 10^{-2}$. Following the common practice employed in [16], we apply a layer-wise learning rate decay of $0.9$ for our VFM encoder. For the experiments conducted on the NYU dataset, we employ RGB images and the HHA encoding of depth images as model input. We conduct all ablation studies on the widely utilized MFNet dataset, with the auxiliary task omitted to ensure training efficiency. The performance improvement (an increase in mIoU by $0.6\%$) between Sect. IV-D and Sect. IV-E demonstrates the effectiveness of the auxiliary task.
IV-D Comparisons with SoTA Scene Parsing Networks
@@
We conduct quantitative comparisons with $12$, $10$, and four SoTA RGB-T scene parsing networks on the MFNet [5], PST900 [1], and KP Day-Night [57] datasets, respectively. The results are detailed in Tables I, III, and II, respectively. Additionally, we present the qualitative comparisons on these three datasets, as shown in Fig. 2, 3, and 4, respectively.
@@
Specifically, our proposed HAPNet achieves the highest mIoU across all datasets, outperforming SoTA methods by $0.1$-$21.8\%$ on the MFNet dataset, $1.0$-$32.0\%$ on the PST900 dataset, and $2.4$-$33.6\%$ on the KP day-night dataset, respectively. These results demonstrate the robustness and effectiveness of HAPNet for RGB-T scene parsing in various scenarios, ranging from complex urban driving scenes to challenging underground scenes. Notably, HAPNet outperforms CRM-RGBTSeg [7], another SoTA Transformer-based approach based on the mask classification paradigm, demonstrating the efficacy of our hybrid, asymmetric architecture in heterogeneous feature fusion. The qualitative results also demonstrate the superior capability of HAPNet in handling various challenging scenes with poor lighting conditions.
@@
In addition, HAPNet demonstrates promising generalizability for RGB-D/HHA scene parsing. As shown in Table IV, our method outperforms all other RGB-T scene parsing networks on the NYU Depth V2 dataset. Specifically, HAPNet achieves an mIoU higher than ECGFNet [49] and RTFNet [4] by $3.5\%$ and $5.9\%$, respectively. Meanwhile, these results suggest that HAPNet still lags behind SoTA methods, developed specifically for RGB-D/HHA data fusion, including Omnivore [51], DFormer-L [40], and OmniVec [50] (mIoU is lower by $2.2$-$5.8\%$). Additionally, it underperforms CMX [8], a SoTA general multi-modal data-fusion method by $1.9\%$ on mIoU. We hypothesize that this performance gap stems from the inherently different data distributions between the thermal and depth/HHA modalities, which limits the applicability of our architecture in the RGB-D/HHA scene parsing domain.
IV-E Ablation Study
@@
We first explore the performance of RGB-T scene parsing with respect to different input data combination strategies. Given the asymmetric two-branch architecture, the RGB-T data can be fed into our HAPNet using nine different strategies to the VFM and CSPD, respectively, as shown in Table V. Among them, strategies E and I can be considered as single-modal versions (only RGB or thermal images are encoded by the network), yielding mIoU scores of only $53.9\%$ and $51.5\%$, respectively. On the other hand, strategies A-D, and F-H, which employ both RGB and thermal images, achieve superior performance, validating the importance of modality complementation. Notably, strategy D, employing RGB images as inputs to the VFM branch while using a combination of RGB-T data in the CSPD, achieves the highest mIoU of $60.9\%$. This result demonstrates that encoding thermal images using the VFM may deteriorate features and degrade the overall performance, which verifies our hypothesis stated in Sect. I.
@@
Our CSPD extracts multi-scale spatial priors from both RGB and thermal images, which are subsequently fused with the VFM context features to implicitly enrich the local semantics. To further explore the most suitable structure for the CSPD, we replace the basic ConvNeXt blocks with other mainstream hierarchical encoder structures, as illustrated in Table VI. Compared to the early ResNet and the latest Swin-Transformer, ConvNeXt exhibits superior performance (yielding an mIoU of $60.9\%$).
@@
Furthermore, we evaluate the effectiveness of GLCA and CCG. As presented in Table VII, removing either of them leads to a significant performance deterioration of HAPNet. The baseline setup, which could only perform cross-modal feature fusion via element-wise summation between spatial priors and global context, exhibits a $3.0\%$ lower mIoU compared to the complete HAPNet.
@@
To validate the effectiveness of our proposed asymmetric architecture, we first construct baseline architectures incorporating symmetric duplex encoders based on BEiTv2 [17] and ConvNeXt [14], respectively. As shown in Table VIII, such baseline architectures fail to achieve performance comparable to that of our developed HAPNet. Specifically, the duplex, symmetric ConvNeXt and BEiTv2 architectures fall short of HAPNet, with a decrease in mIoU by $3.5$-$2.4\%$ and $27.2$-$25.9\%$, respectively. These findings lend further support to our hypothesis stated in Sect. I, suggesting that applying asymmetric architectures to RGB and thermal modalities can play to their own strengths. Importantly, the disappointing performance of the symmetric duplex BEiTv2 encoder suggests that directly applying VFMs for the RGB-T scene parsing task might not be suitable.
@@
Moreover, we also evaluate the compatibility of HAPNet with various SoTA VFMs, including those trained through traditional supervised learning (DeiT [55] and AugReg [56]) as well as the most recent models developed based on self-supervised pre-training (BEiT [16], BEiTv2 [17], and DINOv2 [18]). As shown in Table IX, BEiTv2 achieves the highest mIoU at $60.9\%$, outperforming all other models on the RGB-T scene parsing task. Given BEiTv2’s superior performance, we empirically determine it to be the most suitable VFM for this task, thereby adopting BEiTv2 as the default VFM within HAPNet.
V Conclusion and Future Work
@@
In this article, we revisited the design of heterogeneous feature extraction and fusion strategies. Drawing upon the inherent characteristics of RGB and thermal data, we developed a novel hybrid, asymmetric network to capitalize on their unique strengths, leveraging the powerful VFM for this specific task. We first developed a cross-modal spatial prior descriptor to capture local semantics jointly from RGB-T data. Additionally, we designed a progressive heterogeneous feature integrator, consisting of a global-local context aggregator and a complementary context generator, to fuse heterogeneous features more effectively. We also introduced an auxiliary task to further enhance the local semantics of the fused features, leading to improved overall performance. Extensive experiments demonstrate that HAPNet not only achieves the state-of-the-art performance for RGB-T scene parsing across three public datasets but also shows great potential to generalize well for RGB-HHA scene parsing in indoor scenarios. Despite its superior performance over other existing approaches, the generalizability of HAPNet for RGB-X (where “X” represents, but is not limited to, depth, surface normal, and so forth) can be further improved. Additionally, considering scene parsing is a common functionality embedded in autonomous cars, mobile robots, and drones, the real-time performance of HAPNet is crucial, which is also a direction for possible future work.

### v3

@@ 4.2 Evaluation Metrics
We employ two widely used evaluation metrics: accuracy (Acc) and IoU, to quantify the scene parsing performance with respect to each category. The mean values of these two metrics across all categories (denoted as mIoU and mAcc, respectively) are also computed to quantify the comprehensive network performance.
4.3 Implementation Details
@@
Our model was trained for 200 epochs on an NVIDIA RTX 3090 GPU, utilizing the AdamW optimizer [72] with an initial learning rate of $2\times 10^{-4}$ and a weight decay of $5\times 10^{-2}$. Following the common practice employed in [20], we apply a layer-wise learning rate decay of $0.9$ for our VFM encoder. For the experiments conducted on the NYU dataset, we employ RGB images and the HHA encoding of depth images as model input. We conduct all ablation studies on the widely utilized MFNet dataset, with the auxiliary task omitted. By comparing the results in Tables 1 and 8, we observe that the performance improvement attributed to the auxiliary task is $0.6\%$ in mIoU (calculated as an average across multiple experiments and subject to a fluctuation of approximately $\pm 0.2\%$). These results demonstrate the effectiveness of the auxiliary task.
4.4 Comparisons with SoTA Scene Parsing Networks
@@
We conduct quantitative comparisons with $15$, $10$, and four SoTA RGB-T scene parsing networks on the MFNet [4], PST900 [1], and KP Day-Night [33] datasets, respectively. The results are detailed in Tables 1, 2, and 3, respectively. Additionally, we present the qualitative comparisons on these three datasets, as shown in Fig. 2, 3, and 4, respectively.
@@
Specifically, our proposed HAPNet achieves the highest mIoU across all datasets, outperforming SoTA methods by $0.1$-$21.8\%$ on the MFNet dataset, $1.0$-$32.0\%$ on the PST900 dataset, and $2.4$-$33.6\%$ on the KP day-night dataset, respectively. These results demonstrate the robustness and effectiveness of HAPNet for RGB-T scene parsing in various scenarios, ranging from complex urban driving scenes to challenging underground scenes. Notably, despite performance variations on rare categories (e.g., ”guardrail”) in the MFNet dataset, our approach consistently outperforms CRM-RGBTSeg [7], a state-of-the-art Transformer-based method using the mask classification paradigm. This demonstrates the efficacy of our hybrid asymmetric architecture for heterogeneous feature fusion. Compared to CMNeXt [48], a general-purpose multi-modal data-fusion method that also employs an asymmetric encoder architecture, HAPNet achieves an improvement of $1.6\%$ in mIoU on the MFNet. This performance improvement can be attributed to our method’s utilization of a more powerful VFM and a more advanced multi-scale cross-attention mechanism that integrates global context from the VFM and multi-scale spatial priors provided by a CNN, in contrast to CMNeXt’s reliance on CNN and pooling-based feature fusion strategies. Consequently, HAPNet generates more discriminative fused features for RGB-T scene parsing. The distinct parsing results in Figs. 2, 3, and 4 also demonstrate the superior capability of HAPNet for visual perception under poor illumination conditions.
@@
In addition, HAPNet demonstrates promising generalizability for RGB-D/HHA scene parsing. As shown in Table 4, our method outperforms all other RGB-T scene parsing networks on the NYU Depth V2 dataset. Specifically, HAPNet achieves a $3.5\%$ higher mIoU than ECGFNet [61], and a $5.9\%$ higher mIoU than RTFNet [3]. However, these results suggest that HAPNet still lags behind SoTA methods, developed specifically for RGB-D/HHA data fusion, including Omnivore [65], DFormer-L [51], and OmniVec [64] (mIoU is lower by $2.2$-$5.8\%$). Additionally, it underperforms CMX and CMNeXt [8, 48], two SoTA general-purpose multi-modal scene parsing methods by $1.9\%$ in mIoU. We hypothesize that this performance gap stems from thermal images providing richer complementary features than depth/HHA images. While depth/HHA images primarily convey geometric information, thermal images capture both fine-grained spatial priors and semantic cues through heat signatures (e.g., distinguishing animate from inanimate objects). These characteristics are particularly well-suited to our proposed CSPD and CCG modules. Consequently, these inherent distributional differences between thermal and depth/HHA modalities limit the applicability of our architecture to RGB-D/HHA scene parsing.
Despite achieving SoTA performance, HAPNet demonstrates competitive real-time performance. Our network contains 169.1169.1M parameters and achieves an inference speed of $\sim 23$ frames per second (FPS) when processing images at a resolution of $480\times 480$ pixels on an NVIDIA RTX 4090-D GPU. This performance nearly meets the real-time processing requirements for autonomous driving systems.
4.5 Ablation Study
@@
We first explore the performance of RGB-T scene parsing with respect to different input data combination strategies. Given the asymmetric two-branch architecture, the RGB-T data can be fed into our HAPNet using nine different strategies to the VFM and CSPD, respectively, as shown in Table 5. Among them, strategies E and I can be considered as single-modal versions (only RGB or thermal images are encoded by the network), yielding mIoU scores of only $53.9\%$ and $51.5\%$, respectively. On the other hand, strategies A-D, and F-H, which employ both RGB and thermal images, achieve superior performance, validating the importance of modality complementation. Notably, strategy D, employing RGB images as inputs to the VFM branch while using a combination of RGB-T data in the CSPD, achieves the highest mIoU of $60.9\%$. This result demonstrates that encoding thermal images using the VFM may deteriorate features and degrade the overall performance, which verifies our hypothesis stated in Sect. 1.
@@
Our CSPD extracts multi-scale spatial priors from both RGB and thermal images, which are subsequently fused with the VFM context features to implicitly enrich local semantics. To further explore the most suitable structure for the CSPD, we replace the basic ConvNeXt blocks with other mainstream hierarchical encoder structures. As detailed in Table 6, ConvNeXt demonstrated superior performance, achieving a mIoU of $60.9\%$, outperforming established architectures such as ResNet, Swin-Transformer, and the Mix Vision Transformer (MiT-B4), the latter being a heavy structure previously employed in another RGB-X scene parsing network [48].
@@
Furthermore, we evaluate the effectiveness of GLCA and CCG. As presented in Table 7, removing either of them leads to a significant performance deterioration of HAPNet. The baseline setup, which could only perform cross-modal feature fusion via element-wise summation between spatial priors and global context, exhibits a $3.0\%$ lower mIoU compared to the complete HAPNet.
@@
To validate the effectiveness of our proposed asymmetric architecture, we first construct baseline architectures incorporating symmetric duplex encoders based on BEiTv2 [21] and ConvNeXt [17], respectively. As shown in Table 8, such baseline architectures fail to achieve performance comparable to that of our developed HAPNet. Specifically, the duplex, symmetric ConvNeXt and BEiTv2 architectures fall short of HAPNet, with a decrease in mIoU by $3.5$-$2.4\%$ and $27.2$-$25.9\%$, respectively. These findings lend further support to our hypothesis stated in Sect. 1, suggesting that applying asymmetric architectures to RGB and thermal modalities can play to their own strengths. Importantly, the disappointing performance of the symmetric duplex BEiTv2 encoder suggests that directly applying VFMs for the RGB-T scene parsing task might not be suitable. We further compare the encoder of HAPNet with another best-performing asymmetric encoder proposed in [48]. In order to fairly compare the performance of the two encoder architectures, we combined CMNeXt’s encoder with HAPNet’s decoder, abbreviated as HAPNet CMNeXt-B4, as shown in Table 8, and the mIoU exhibits a decrease of $1.6\%$, further underscoring the superiority of our encoder design compared to other architectures for RGB-T scene parsing.
@@
Moreover, we also evaluate the compatibility of HAPNet with various SoTA VFMs, including those trained through traditional supervised learning (DeiT [69] and AugReg [70]) as well as the most recent models developed based on self-supervised pre-training (BEiT [20], BEiTv2 [21], and DINOv2 [22]). As shown in Table 9, BEiTv2 achieves the highest mIoU of $60.9\%$, outperforming all other models on the RGB-T scene parsing task. Given BEiTv2’s superior performance, we empirically determine it to be the most suitable VFM for this task, thereby adopting BEiTv2 as the default VFM within HAPNet.
5 Conclusion and Future Work
@@
In this article, we revisited the design of heterogeneous feature extraction and fusion strategies. Drawing upon the inherent characteristics of RGB and thermal data, we developed a novel hybrid, asymmetric network to capitalize on their unique strengths, leveraging the powerful VFM for this specific task. We first developed a cross-modal spatial prior descriptor to capture local semantics jointly from RGB-T data. Additionally, we designed a progressive heterogeneous feature integrator, consisting of a global-local context aggregator and a complementary context generator, to fuse heterogeneous features more effectively. We also introduced an auxiliary task to further enhance the local semantics of fused features, leading to improved overall performance. Extensive experiments demonstrate that HAPNet not only achieves state-of-the-art performance for RGB-T scene parsing across three public datasets but also shows great potential to generalize well for RGB-HHA scene parsing in indoor scenarios. Despite its superior performance over other existing approaches, the generalizability of HAPNet for RGB-X (where “X” represents, but is not limited to, depth, surface normal, and so forth) can be further improved. Additionally, considering scene parsing is a common functionality embedded in autonomous cars, mobile robots, and drones, the real-time performance of HAPNet is crucial, which is also a direction for possible future work.
6 Acknowledgements
This research was supported by the National Natural Science Foundation of China under Grant 62473288, the Fundamental Research Funds for the Central Universities, and the Xiaomi Young Talents Program.
Table 10: Backbone architectures, model complexities (in total parameters, Params), and inference speeds (in frames per second, FPS) for all evaluated networks, with all performance benchmarks conducted on an NVIDIA RTX 3090 GPU. “-“ denotes that the corresponding results are unavailable.
Method Backbone Params (M) FPS Method Backbone Params (M) FPS
MFNet  [4] Mini-Inception 0.72 194.0 CMX  [8] MiT-B4 139.9 11.72
RTFNet  [3] ResNet-152 254.51 26.18 CMNeXt  [48] MiT-B4 119.6 12.6
FuseSeg  [29] DenseNet-161 141.52 23.85 CRM-RGBTSeg  [7] Swin-B 193.0 -
EGFNet  [60] ResNet-152 62.82 8.58 GMNet  [9] ResNet-50 149.8 15.0
ABMDRNet  [5] ResNet-50 64.60 43.46 Omnivore  [65] Swin-B 95.7 -
ECGFNet  [61] MobileNetV2 26.06 38.37 TokenFusion  [66] MiT-B3 45.9 -
FEANet  [25] Modified ResNet 255.21 21.13 DFormer-L  [51] DFormer-L 39.0 -
SFAF-MA  [62] ResNet-152 156.0 21.8 LLE-Seg  [63] ConvNeXt V2 - -
ABMDRNet+  [24] ResNet-50 64.60 40.14 AsymFormer  [67] Modified MiT-B0 33.0 -
CAINet  [10] MobileNetV2 12.16 61.44 SA-Gate  [68] ResNet-50 110.85 18.36
EAEFNet  [2] ResNet-152 200.4 21.0 OmniVec  [64] ViT varients 95.7 -
Appendix A Mathematical Details
This section provides details on the mathematical derivations of the multi-head self-attention (MHSA) [73] and multi-scale deformable cross-attention (MSDA) [31], which are integral components of our proposed HAPNet. MHSA is employed within the vision foundation model, while MSDA is utilized in the progressive heterogeneous feature aggregators.
A.1 Multi-head attention
Let $\boldsymbol{X}_{q}\in\mathbb{R}^{N_{q}\times C}$ be the query, and let $\boldsymbol{X}_{f}\in\mathbb{R}^{N_{f}\times C}$ be the key and value, where $N_{q}$ and $N_{f}$ are the respective sequence lengths. The multi-head attention (MHA), including both MHSA and multi-head cross-attention, can be formulated as follows:
$\boldsymbol{Q}^{i},\boldsymbol{K}^{i},\boldsymbol{V}^{i}=\boldsymbol{X}_{q}\boldsymbol{U}_{q}^{i},\,\boldsymbol{X}_{f}\boldsymbol{U}_{k}^{i},\,\boldsymbol{X}_{f}\boldsymbol{U}_{v}^{i},\,\,\,i\in[1,n]\cap\mathbb{Z},$ (7)
$\boldsymbol{A}^{i}=\operatorname{Softmax}(\frac{\boldsymbol{Q}^{i}(\boldsymbol{K}^{i})^{\top}}{\sqrt{C_{v}}})$ (8)
$\text{MHA}(\boldsymbol{X}_{q},\boldsymbol{X}_{f})=\operatorname{Concat}([\boldsymbol{A}^{1}\boldsymbol{V}^{1};\dots;\boldsymbol{A}^{n}\boldsymbol{V}^{n}])\boldsymbol{U}_{m},$ (9)
where $\boldsymbol{U}_{q}^{i}\in\mathbb{R}^{C\times C_{v}}$, $\boldsymbol{U}_{k}^{i}\in\mathbb{R}^{C\times C_{v}}$ and $\boldsymbol{U}_{v}^{i}\in\mathbb{R}^{C\times C_{v}}$ denotes the linear projections of the query, key, and value in the $i$-th attention head, respectively, $\boldsymbol{A}^{i}$ stores attention weights, and $\boldsymbol{U}_{m}\in\mathbb{R}^{n\cdot C_{v}\times C}$ represents the linear projection of $n$ attention operations ($C_{v}=\frac{C}{n}$ by default). $\operatorname{Concat}([\cdot;\dots;\cdot])$ denotes the feature concatenation operation performed on the results from $n$ attention operations.
A.2 Deformable attention
Compared to standard MHA, the deformable attention (DA) only attends a small set of points from the feature map as the reference points instead of the whole feature map, which reduces the computation complexity while maintaining competitive performance. The sampling process of DA can be achieved by predicting the sampling offsets and attention weights $\boldsymbol{A}^{i}$ through linear projections of the query $\boldsymbol{X}_{q}$. Let $\boldsymbol{p}_{q}=(h,w)$ index a 2-D reference point on the resized $\boldsymbol{X}_{q}$, where $\boldsymbol{X}_{q}$ has been restored to its original 2-D spatial dimensions. At this reference point, $\boldsymbol{X}_{q}(\boldsymbol{p}_{q})\in\mathbb{R}^{1\times C}$ represents the feature vector corresponding to position $\boldsymbol{p}_{q}$. The DA operation can be formulated as:
$\operatorname{DA}^{i}(\boldsymbol{X}_{q}(\boldsymbol{p}_{q}),\boldsymbol{X}_{f})=\sum_{k=1}^{K}\boldsymbol{A}^{i}_{k,\boldsymbol{p}_{q}}\cdot\boldsymbol{X}_{f}(\boldsymbol{p}_{q}+\Delta^{i}_{k,\boldsymbol{p}_{q}})\boldsymbol{U}_{v}^{i},$ (10)
where $k$ indexes the sampled keys, and $K$ is the total number of sampled keys. $\Delta^{i}_{k,\boldsymbol{p}_{q}}$ and $\boldsymbol{A}^{i}_{k,\boldsymbol{p}_{q}}$ denote the sampling offsets and attention weights of the $k$-th sampling point on $\boldsymbol{p}_{q}$ in the $i$-th attention head, respectively. The attention weights $\boldsymbol{A}^{i}_{k,\boldsymbol{p}_{q}}$ are normalized to satisfy $\sum_{k=1}^{K}\boldsymbol{A}^{i}_{k,\boldsymbol{p}_{q}}=1$.
Let $\mathcal{X}_{f}=\{\boldsymbol{X}^{l}_{f}\}_{l=1}^{L}$ be the multi-scale feature maps, where $\boldsymbol{X}^{l}_{f}\in\mathbb{R}^{H_{l}\times W_{l}\times C}$, and let $\boldsymbol{\hat{p}}_{q}\in[0,1]^{2}$ be the normalized coordinate of the reference point for the query $\boldsymbol{X}_{q}$. The $\operatorname{DA}$ process can be extended to its MSDA form as follows:
$\operatorname{MSDA}^{i}(\boldsymbol{X}_{q}(\boldsymbol{\hat{p}}_{q}),\mathcal{X}_{f})=\sum_{l=1}^{L}\sum_{k=1}^{K}\boldsymbol{A}^{i,l}_{k,\boldsymbol{p}_{q}}\cdot\boldsymbol{X}^{l}_{f}(\phi_{l}(\boldsymbol{\hat{p}}_{q})+\Delta^{i,l}_{k,\boldsymbol{p}_{q}})\boldsymbol{U}_{v}^{i},$ (11)
where $l$ indexes the feature map level, and $k$ indexes the sampling point. $\Delta^{i,l}_{k,\boldsymbol{p}_{q}}$ and $\boldsymbol{A}^{i,l}_{k,\boldsymbol{p}_{q}}$ denote the corresponding sampling offsets and attention weights, respectively, for the $k$-th sampling point in the $l$-th feature level and the $i$-th attention head. The attention weights $\boldsymbol{A}^{i,l}_{k,\boldsymbol{p}_{q}}$ are normalized to satisfy $\sum_{l=1}^{L}\sum_{k=1}^{K}\boldsymbol{A}^{i,l}_{k,\boldsymbol{p}_{q}}=1$. The transformation $\phi_{l}(\cdot)$ rescales the $\boldsymbol{\hat{p}}_{q}$ in (11) to the original spatial size of the $l$-th feature level.
Appendix B Supplementary Content for Tables
This section provides the detailed backbone configurations, model complexities, and inference speeds for all methods compared with HAPNet in this study, as shown in Table 10.